A helium balloon occupies 112mL at 90.0 kPa and 34 degrees Celsius. What volume would it occupy at standard temperature and pressure (101.3 kPa & 0.00 degrees Celsius)
1
Expert's answer
2020-05-04T01:53:06-0400

Use the combined gas law.

P1V1/T1 = P2V2/T2

90*0.112/307 = 101.3*V2/273

0.0328 = 0.3711*V2

V2 = 0.088L